Judy and Mary started walking toward each other from two different locations. They started at the same time. Judsy walks 2 miles per hour and Mary walks 2.5 miles per hour. By the time Judy reached mid point, they were 3 miles apart. How far apart were they when they started?

Answer by stanbon(75887)

Judy and Mary started walking toward each other from two different locations. They started at the same time. Judy walks 2 miles per hour and Mary walks 2.5 miles per hour. By the time Judy reached mid point, they were 3 miles apart. How far apart were they when they started?
------------------------------------------------
Judy DATA:
rate = 2 mph ; time = x hrs ; distance = 2x miles
--------------
Mary DATA:
rate = 2.5 mph ; time = 2.5 hrs ; distance = 2.5x miles
---------------
Equation:
distance - distance = 3 miles
2.5x-2x = 3 miles
0.5x = 3
x = 6 hrs
================
Judy walked 2*6 = 12 miles
Mary walked 2.5*6 = 15 miles
----
They started 2*12 = 24 miles apart.
======================================
Cheers,
Stan